# Diffusion-based Generative Multicasting with Intent-aware Semantic Decomposition

### v1

@@ Abstract
Generative diffusion models (GDMs) have recently shown great success in synthesizing multimedia signals with high perceptual quality enabling highly efficient semantic communications in future wireless networks. In this paper, we develop an intent-aware generative semantic multicasting framework utilizing pre-trained diffusion models. In the proposed framework, the transmitter decomposes the source signal to multiple semantic classes based on the multi-user intent, i.e. each user is assumed to be interested in details of only a subset of the semantic classes. The transmitter then sends to each user only its intended classes, and multicasts a highly compressed semantic map to all users over shared wireless resources that allows them to locally synthesize the other classes, i.e. non-intended classes, utilizing pre-trained diffusion models. The signal retrieved at each user is thereby partially reconstructed and partially synthesized utilizing the received semantic map. This improves utilization of the wireless resources, with better preserving privacy of the non-intended classes. We design a communication/computation-aware scheme for per-class adaptation of the communication parameters, such as the transmission power and compression rate to minimize the total latency of retrieving signals at multiple receivers, tailored to the prevailing channel conditions as well as the users’ reconstruction/synthesis distortion/perception requirements. The simulation results demonstrate significantly reduced per-user latency compared with non-generative and intent-unaware multicasting benchmarks while maintaining high perceptual quality of the signals retrieved at the users. For a typical setup of multicasting street scene images to 10 users, our proposed framework achieves a $15.4\%$ reduction in per-user latency at a fixed power budget, or equivalently $50\%$ reduction in the transmission power required to achieve a fixed per-user latency, compared with non-generative multicasting.
@@ I Introduction
Semantic Communication (Semcom) is envisioned to play a crucial role in future wireless networks, specifically in emerging applications where communication of large multimodal signals with stringent latency and reliability constraints is required, e.g. the wireless metaverse and digital twins [1, 2, 3], extended/mixed reality (XR/MR) [4, 5], and holographic teleportation and the internet of senses [6, 7]. In many such applications, multicasting of multimodal signals is required, where the wireless and network resource utilization can be significantly improved by transmitting the same desired content to multiple users over shared resources. For example in the wireless metaverse, multicasting is used to communicate updates, states, or events of the virtual world to multiple users simultaneously in real-time to ensure all users remain synchronized. Another example is XR/MR streaming, where the same content may be distributed to multiple users simultaneously. In these applications, Semcom can further reduce the wireless and network resource utilization by extracting and communicating only the semantic content of interest based on the users’ intent [8, 9, 10, 11, 12].
Recently, the Generative AI (GenAI) models have shown great success in developing efficient low bitrate Generative Semantic Communication (Gen Semcom) systems [13, 14, 15, 16]. Generative models are capable of learning the general distribution of natural signals during training and synthesizing new samples with high perceptual quality at inference time. The generation process can be guided by multi-modal prompts and conditioning signals to produce high quality outputs with a desired semantic content. The emerging Generative Foundation Models and Multimodal Large Language Models (MLLMs), e.g. Sora [17], Lumiere [18], and DALL.E [19], etc., provide ample opportunities to develop efficient and universal generative Semcom frameworks [20, 21, 22, 23]. As these models possess a vast general knowledge captured via intensive pre-training on huge amount of data, they alleviate the need for a shared Knowledge Base (KB) between the semantic transmitter and receiver, thereby reducing the corresponding knowledge sharing overheads. This vast general knowledge also makes the generative Semcom framework applicable to various datasets and tasks thereby achieving universality. Moreover, generative Semcom offers an inherent advantage in privacy preservation, as only the intended features of the original signal are communicated, and the non-intended details are generated locally from random noise inputs at the receiver utilizing the generative model. Such details are thereby not precisely reproducible and are kept private from the receiver, without requiring separate anonymization techniques
@@
The GenAI models are trained to maximize the perceptual quality of the synthesized signal, and the theoretical limits of generative Semcom are governed by the rate-distortion-perception theory [24, 25, 26, 27]. This theory explores the threefold trade-off between rate, distortion, and the perceptual quality of the synthesized signal. While the rate-distortion-perception function is analytically derived only for a few source distributions, it is generally estimated empirically for natural signals, e.g. images, audio/video, point cloud, etc., by training deep source encoders with a perceptual loss function, e.g. Wasserstein distance, on large datasets [24, 25]. In generative Semcom, the transmitter extracts the intended semantics, e.g. in form of textual prompts [28, 29], compressed embeddings [30, 31], semantic/edge map [15, 21, 32] etc., which are then transmitted over the channel. The receiver uses these semantics to guide a generative model, synthesizing a signal that is semantically consistent and highly realistic. Despite the above existing works, multiuser generative semantic communications, specifically in the multicasting setup, and its scalability to increasing number of users under latency constraints remains less studied.
@@
In this paper, we develop a generative semantic multicasting framework with adaptive intent-aware resource allocation utilizing pre-trained diffusion models. We claim two key benefits for the proposed framework: Firstly, generative multicasting in our Semcom framework allows preserving privacy of the non-intended signals, while improving utilization of the wireless and network resources compared with non-generative and intent-unaware multicasting benchmarks. This also reduces the total latency, while maintaining the reconstruction/synthesis distortion/perception quality. Secondly, the adoption of pre-trained diffusion models allows a source-channel separation-based Semcom architecture, thereby alleviating the need for end-to-end joint training of the transmitter and receiver, which is required in many existing Semcom frameworks [8, 9, 10, 11]. Such a separation-based architecture offers improved adaptability to varying channel conditions, improved scalability to increasing number of users, and better compatibility with the existing design of wireless communication networks. The contributions of this work are three-fold:
@@
  • We develop a generative multicasting framework where a semantic decomposition scheme splits the source signal into multiple semantic classes at the transmitter based on the multi-user intent, i.e. each user is assumed to need only a subset of the semantic classes based on its communication intent. The transmitter sends to each user only its intended classes over orthogonal wireless resources, and multicasts to all users a highly compressed semantic map over shared resources. Each user then locally synthesizes the non-intended classes using a pre-trained generative diffusion model.
@@
  • For a typical image synthesis task, we train a generative diffusion model with classifier-free semantic guidance to derive estimates of the corresponding rate-distortion/perception curves achieved by the model through pre-training. We use these curves to determine the communication parameters, i.e. the transmission power and compression rate in our proposed framework.
@@
  • We design a communication/computation-aware scheme for per-class adaptation of the communication parameters to minimize the total latency for varying channel conditions and users’ reconstruction/synthesis distortion/perception requirements. The signal retrieved at each user is partially reconstructed and partially synthesized utilizing the pre-trained diffusion model, and the communication parameters are optimized for latency-aware synchronisation of the synthesized and reconstructed datastreams based on the prevailing channel conditions and multi-user intent.
@@
The rest of this paper is organized as follows. In Section II we provide the related works. In Section III, we present our proposed diffusion-based generative multicasting framework and the corresponding intent and latency-aware adaptive semantic communication scheme. Simulation results are provided in Section IV, and we conclude the paper in Section V.
@@
II Related Works
In this section, we introduce the related works on denoising diffusion models and semantic guidance for visual synthesis.
II-A Denoising Diffusion Models
Denoising diffusion models are a class of generative models that learn to generate samples of natural signals (e.g. image, video, etc.) by reversing a diffusion process, in which the signal is progressively noised until it becomes indistinguishable from random noise. Diffusion models are parameterized Markov chains that learn to generate samples matching the data distribution by training the model parameters to optimize the lower variational bound on the likelihood function [33, 34, 35]. To generate samples the model starts with random noise input and gradually refines it in iterations, to produce a clear structured output. Diffusion models have recently demonstrated state of the art performance for generating multimedia content such as images [36, 37], audio/video signals [38, 39, 40], and 3D shapes [41, 42]. With the development of novel diffusion models for various generation tasks, many free-to-public demos have been released, equipped with models pretrained on extra-large-scale data that demonstrate remarkable capabilities to generalize on different tasks. Open source availability crucially boosts the adoption of diffusion models, allowing them to be incorporated into other novel frameworks and a plethora of applications. We can identify two key factors for the recent success and popularity of diffusion models. Firstly, the generation ability of diffusion models crucially outperforms the capabilities of other models, such as Generative Adversarial Networks (GANs), flow-based models, or Variational Autoencoders (VAEs) [35]. Secondly, the training and sampling process of diffusion models is far more stable than the generation process of GANs, making them a more reliable method [43]. These factors favor the adoption of diffusion models in generative Semcom, where models’ ability to synthesize high quality content from extremely compressed information with affordable computational complexity is crucial.
II-B Semantic Guidance for Visual Synthesis
Semantic visual synthesis models transform semantic maps/layouts into diverse realistic images/videos/3D media. These models are conditioned/guided by the semantic label maps to generate semantically consistent outputs. State-of-the-art semantic synthesis models are mostly based on diffusion models [44, 36, 37] or GANs [45, 46, 47, 48]. Some of the GAN-based models include Pix2PixHD [45], SPADE [46], CLADE [47] for images, and SVS-GAN [48] for video synthesis. More recent works are based on diffusion models. Latent Diffusion Models (LDMs) [36] apply diffusion models in the lower-dimensional latent space of pretrained autoencoders to make high-resolution image synthesis more computationally efficient. Cross-attention modules are used in LDMs to inject spatial conditioning information such as semantic/edge/depth maps into the latent diffusion process. In ControlNet [37], a neural network architecture is proposed that can add conditional guidance to large pretrained generative diffusion models with little further fine-tuning. In this work, we adopt the Semantic Diffusion Model (SDM) [44], which introduces classifier-free semantic guidance to enhance the quality of generation by modulating the influence of the conditioning signal during diffusion, without requiring an explicit classifier. This approach combines guided and unguided samples for training to balance creativity and fidelity of the model. SDM is trained using a combination of a simple denoising loss, where the model learns to predict the noise introduced during the diffusion process, and a variational loss that minimizes the KL divergence between the predicted and true posterior distributions. This two-pronged approach ensures that the model generates images that are both realistic and closely aligned with the input semantic maps.
@@ II-B Semantic Guidance for Visual Synthesis
III Intent-aware Generative Semantic Multicasting
We consider a transmitter serving $K$ users in a multicasting setup as depicted in Fig. 1. The source signal to be transmitted to the users contains $L$ semantic classes. For example, assuming the source signal is images captured by a street scene camera, the semantic classes include “car”, “person”, “building”, “sky”, etc. Denoting the source signal by ${\bf X}$, we have ${\bf X}=\bigcup_{l=1}^{L}{\bf X}_{l}$, where ${\bf X}_{l}$ represents the $l$’th class. Each user is only interested in a specific subset of the classes based on its goal or intent. For example, a user carrying out “traffic surveillance” based on the received signal is interested in the “car” class but not “sky”. This user may need details of the car class, e.g. the cars’ make and model, as well as their color and number plate. A user carrying out “weather monitoring” is only interested in details of the “sky” class, and another user carrying out “Path Planning” for Intelligent Transportation Systems (ITS) will be interested in the “street signs” and the “road” classes. Hence, we define the multi-user intent matrix ${\bf I}_{K\times L}$, where $[{\bf I}]_{kl}=1$ if user $k$ is interested in class $l$, otherwise $[{\bf I}]_{kl}=0$. The intended portion of the source signal for use $k$, i.e. the semantic classes each user is interested in, is thereby written as ${\bf X}^{(k)}_{I}=\{\bigcup_{l}{\bf X}_{l}|[{\bf I}]_{kl}=1\}$. Similarly, the non-intended portion of the source signal for use $k$ is thereby written as ${\bf X}^{(k)}_{N}=\{\bigcup_{l}{\bf X}_{l}|[{\bf I}]_{kl}=0\}$.
@@ III Intent-aware Generative Semantic Multicasting
III-A Multi-class Semantic Decomposition for Mixed Reconstruction/Synthesis with Semantic Diffusion Models
III-A1 Multi-class Semantic Decomposition with DDRNets
@@
III-A2 Mixed Reconstruction/Synthesis with Semantic Diffusion Models
At the receiver, each user receives two datastreams, one containing its intended classes which is decompressed to reconstruct its intended portion of the source signal, and the other containing the semantic map ${\bf M}$ which is fed into a pre-trained generative diffusion model for high-fidelity synthesis of its non-intended portion. The signal retrieved at user $k$ is thereby written as ${\bf Y}^{(k)}={\bf Y}^{(k)}_{r}\bigcup{\bf Y}^{(k)}_{s}$, where ${\bf Y}^{(k)}_{r}$ is the reconstructed and ${\bf Y}^{(k)}_{s}$ is the synthesized portion. The proposed mixed reconstruction/synthesis approach improves the communication privacy as it avoids revealing details of the non-intended classes to users that do not need that information. Any state-of-the-art generative models with semantic guidance can be used to locally synthesize the non-intended signal portions ${\bf Y}^{(k)}_{s}$ at the users. Here we pre-train GDMs with classifier-free guidance for high quality/fidelity image synthesis from the semantic map at each user. Dropping the user index $k$ for brevity, the forward (diffusion) process is defined via
@@ III-A2 Mixed Reconstruction/Synthesis with Semantic Diffusion Models
where ${\bf Y}_{t}$ is the signal at the $t$’th diffusion step, $\rho^{\mathrm{Diff}}({\bf Y}_{t}\mid{\bf Y}_{t-1})$ is the diffusion distribution, and $\beta_{t}\in(0,1)$ is the trainable variance schedule controlling the amount of noise added at each step. To synthesize ${\bf Y}_{s}={\bf Y}_{0}$, the reverse (synthesis) process learns to remove the noise iteratively, transforming the noisy input ${\bf Y}_{\tau}$ to the synthesized image ${\bf Y}_{0}$, using the additional conditioning information from the received semantic map ${\bf\hat{M}}$. This is modeled by a parameterized Markov chain
$\rho^{\mathrm{Syn}}_{\theta}({\bf Y}_{t-1}\mid{\bf Y}_{t},{\bf\hat{M}})$ $=\mathcal{N}({\bf Y}_{t-1};\mu_{\theta}({\bf Y}_{t},{\bf\hat{M}},t),\Sigma_{\theta}({\bf Y}_{t},{\bf\hat{M}},t)),$ (2)
$\rho^{\mathrm{Syn}}_{\theta}({\bf Y}_{0:\tau}\mid{\bf\hat{M}})$ $=p({\bf Y}_{\tau})\prod_{t=1}^{\tau}p_{\theta}({\bf Y}_{t-1}\mid{\bf Y}_{t},{\bf\hat{M}}),$
where $\rho^{\mathrm{Syn}}_{\theta}({\bf Y}_{t-1}\mid{\bf Y}_{t},{\bf\hat{M}})$ is a Gaussian distribution with learned parameters $\mu_{\theta}({\bf Y}_{t},{\bf\hat{M}},t)$ and $\Sigma_{\theta}({\bf Y}_{t},{\bf\hat{M}},t)$, in which $\mu_{\theta}$ and $\Sigma_{\theta}$ are learned functions that depend on the noisy image ${\bf Y}_{t}$, the semantic map ${\bf\hat{M}}$, and the current timestep $t$. To learn these parameters, we use a U-Net architecture where the noisy image is passed through the encoder, consisting of a series of downsampling ResBlocks and attention layers. These blocks operate at various resolutions to capture spatial features at multiple scales. The semantic map information is injected into the decoder using SPADE architectures [46]. At each layer of the decoder, SPADE scales the features according to the semantic map, embedding semantic information in a spatially adaptive manner. This allows the network to maintain a strong correlation between the generated image and the input semantic map.
The goal of training is to minimize the variational lower bound (VLB) of the negative log-likelihood. This can be broken down into a denoising objective by optimizing a noise prediction model. Specifically, the model learns to predict the noise $\epsilon$ at each step, and the loss is formulated as
@@
where $\epsilon$ is the true noise sampled from a Gaussian distribution, $\epsilon_{\theta}$ is the noise predicted by the model at time step $t$, given the noisy image ${\bf Y}_{t}$, the semantic map ${\bf\hat{M}}$, and the time step $t$. During inference, the diffusion process iteratively denoises the noisy image input ${\bf Y}_{\tau}$ to generate the final image ${\bf Y}_{0}$. At each step $t$, the model predicts the noise $\epsilon_{\theta}({\bf Y}_{t},{\bf\hat{M}},t)$ and updates the image as
${\bf Y}_{t-1}=\frac{1}{\sqrt{1-\beta_{t}}}\left({\bf Y}_{t}-\frac{\beta_{t}}{\sqrt{1-\alpha_{t}}}\epsilon_{\theta}({\bf Y}_{t},{\bf\hat{M}},t)\right)+\sigma_{t}{\bf N}$
@@
III-B Intent-Aware Generative Multicasting

### v2

# Communicate Less, Synthesize the Rest: Latency-aware Intent-based Generative Semantic Multicasting with Diffusion Models

@@ Abstract
Generative diffusion models (GDMs) have recently shown great success in synthesizing multimedia signals with high perceptual quality enabling highly efficient semantic communications in future wireless networks. In this paper, we develop an intent-aware generative semantic multicasting framework utilizing pre-trained diffusion models. In the proposed framework, the transmitter decomposes the source signal to multiple semantic classes based on the multi-user intent, i.e. each user is assumed to be interested in details of only a subset of the semantic classes. To better utilize the wireless resources, the transmitter sends to each user only its intended classes, and multicasts a highly compressed semantic map to all users over shared wireless resources that allows them to locally synthesize the other classes, i.e. non-intended classes, utilizing pre-trained diffusion models. The signal retrieved at each user is thereby partially reconstructed and partially synthesized utilizing the received semantic map. We design a communication/computation-aware scheme for per-class adaptation of the communication parameters, such as the transmission power and compression rate to minimize the total latency of retrieving signals at multiple receivers, tailored to the prevailing channel conditions as well as the users’ reconstruction/synthesis distortion/perception requirements. The simulation results demonstrate significantly reduced per-user latency compared with non-generative and intent-unaware multicasting benchmarks while maintaining high perceptual quality of the signals retrieved at the users. For a typical setup of multicasting street scene images to 10 users, our proposed framework achieves a $15.4\%$ reduction in per-user latency at a fixed power budget, or equivalently $50\%$ reduction in the transmission power required to achieve a fixed per-user latency, compared with non-generative multicasting.
@@ I Introduction
Semantic Communication (Semcom) is envisioned to play a crucial role in future wireless networks, specifically in emerging applications where communication of large multimodal signals with stringent latency and reliability constraints is required to reduce the wireless/network resources used, e.g. the wireless metaverse and digital twins, extended/mixed reality (XR/MR), and holographic teleportation and the internet of senses. In many such applications, different receivers may be interested in different parts of the signal according to their communication intent. Considering the limited wireless/network resources, the traditional multicasting schemes that transmit the same content to all users irrespective of their intent, perform inefficiently in such scenarios, thereby necessitating development of new intent-aware SemCom schemes for multicasting with intent heterogeneity. As an example, consider an XR/MR streaming scenario where the content needs to be simultaneously distributed to multiple users with different intents, or in the wireless metaverse, intent-aware multicasting is required to communicate the intended updates, states, or events of the virtual world to multiple users simultaneously in real-time to ensure all users remain synchronized. In many similar applications, multicasting to multiple users with different semantic interests is essential. Despite this need, the existing works in SemCom focus mostly on single-user point-to-point scenarios, and semantic multicasting remains less studied. In this work, we propose generative semantic multicasting, where intent-aware decomposition of the source signal is carried out at the transmitter to allow communicating to each user only its intended part of the signal and synthesizing the rest locally with diffusion models, thereby avoiding unnecessary use of the wireless/network resources for transmission of the non-intended portions of the signal.
@@
Recently, the Generative AI (GenAI) models have shown great success in developing efficient low bitrate Generative Semantic Communication (Gen Semcom) systems [2, 3, 4, 5]. Generative models are capable of learning the general distribution of natural signals during training and synthesizing new samples with high perceptual quality at inference time. The generation process can be guided by multi-modal prompts and conditioning signals to produce high quality outputs with a desired semantic content. The emerging Generative Foundation Models and Multimodal Large Language Models (MLLMs), e.g. Sora, Lumiere, and DALL.E, etc., provide ample opportunities to develop efficient generative Semcom frameworks [6, 7, 8, 9, 10]. As these models possess a vast general knowledge captured via intensive pre-training on huge amount of data, they alleviate the need for a shared Knowledge Base (KB) between the semantic transmitter and receiver, thereby reducing the corresponding knowledge sharing overheads. This vast general knowledge also makes the generative Semcom framework applicable to various datasets and tasks thereby achieving universality. The GenAI models are trained to maximize the perceptual quality of the synthesized signal, and the theoretical limits of generative Semcom are governed by the rate-distortion-perception theory [11, 12]. This theory explores the threefold trade-off between rate, distortion, and the perceptual quality of the synthesized signal. While the rate-distortion-perception function is analytically derived only for a few source distributions, it is generally estimated empirically for natural signals, e.g. images, audio/video, point cloud, etc., by training deep source encoders with a perceptual loss function, e.g. Wasserstein distance, on large datasets. In generative Semcom, the transmitter extracts the intended semantics, e.g. in form of textual prompts [13, 14, 15], compressed embeddings [16, 17], semantic map [18, 19], edge map [2, 7], etc., which are then transmitted over the channel. The receiver uses these semantics to guide a generative model, synthesizing a signal that is semantically consistent and highly realistic. Despite the above works, multiuser generative semantic communications, specifically in the multicasting setup, and its scalability to increased number of users, considering the communication/computation latency remains less studied.
[20] [21] [22] [23] This Work
JSCC or SSCC? SSCC JSCC SSCC JSCC SSCC
E2E Training Needed? Yes Yes Yes Yes No (Pre-Training)
Gen or Non-Gen? Non-Gen Non-Gen Non-Gen Gen (MAE) Gen (Diffusion)
Intent/Task Heterogeneity? Yes Yes No Yes Yes
Comm/Comp. Latency Aware? No No No No Yes
TABLE I: Key Differences Between this Work and Semantic Multicasting/Broadcasting Literature.
@@
In this paper, we develop an scalable latency-aware generative semantic multicasting framework with adaptive intent-based resource allocation utilizing pre-trained diffusion models. We claim three key benefits for the proposed framework: Firstly, generative multicasting in our Semcom framework improves utilization of the wireless and network resources compared with non-generative and intent-unaware multicasting benchmarks. This also reduces the total latency while maintaining the reconstruction/synthesis distortion/perception quality. Secondly, the adoption of pre-trained models allows a source-channel separation-based Semcom architecture, thereby alleviating the need for end-to-end joint training of the transmitter and receiver, which is required in many existing Semcom frameworks [24, 25, 26, 27]. Such a separation-based architecture offers improved adaptability to varying channel conditions, improved scalability to increase the number of users, and better compatibility with the existing design of wireless communication networks. Finally, diffusion models enable synthesizing photorealistic outputs with better diversity, realism, and fine-grained details. TABLE I provides an overview of the major differences between this work and the key existing works on semantic multicasting/broadcasting [20, 21, 22, 23]. To the best of our knowledge, this work is the first of its kind to design joint communication/computation latency-aware intent-based generative semantic multicasting with pre-trained diffusion models, based on source-channel coding separation to achieve scalable and adaptable semantic multicasting compatible with the existing wireless networks, and derive the corresponding rate-distortion/perception tradeoffs. The contributions of this work are thereby three-fold:
@@
  • To achieve intent-awareness, we propose a semantic decomposition scheme which splits the source signal into multiple sub-signal classes at the transmitter based on the multi-user intent feedback, i.e. each user is assumed to need only a subset of the semantic classes based on its communication intent. We develop a mixed reconstruction/synthesis scheme with semantic diffusion models, where users partially reconstruct and partially synthesize the signal from a highly compressed semantic map, i.e. the intended sub-signal classes are reconstructed and the non-intended sub-signals classes are locally synthesized by the pre-trained generative diffusion model. The transmitter broadcasts the semantic map to all users over shared wireless resources, thereby utilizing orthogonal resources only to transmit the sub-signal classes intended for each user, which further saves the wireless resources.
@@
  • To achieve scalability and adaptability, different from majority prior SemCom research, we follow a septate source-channel coding scheme and pre-train a generative diffusion model with classifier-free semantic guidance. Accordingly, we derive the achievable rate-distortion/perception curves for our proposed mixed reconstruction/synthesis scheme. We then use these curves to determine and adapt the communication parameters, i.e. the transmission power and compression rate in our proposed framework on the go, without requiring end-to-end retraining. We show efficient scalability and adaptability of our proposed framework with extensive simulations.
@@
  • To achieve latency-awareness, we design a scheme for per-class adaptation of the communication parameters to minimize the total latency considering both the communication and computation time. As the signal retrieved at each user is partially reconstructed and partially synthesized, we formulate optimization of the communication parameters for latency-aware synchronisation of the synthesized and reconstructed datastreams, based on the prevailing channel conditions, users’ intents and their reconstruction/synthesis distortion/perception requirements. We propose a computationally efficient sequential quadratic programming (SQP)-based algorithm for solving the resulting optimization problem, demonstrating s a 15.4% reduction in per-user latency at a fixed power budget, or equivalently 50% reduction in the transmission power required to achieve a fixed per-user latency compared with intent-unaware and non-generative multicasting, for a typical image semantic multicasting setup of street scene images with 10 users.
@@
The rest of this paper is organized as follows. In Section II, we present our proposed diffusion-based generative semantic communication framework and the corresponding intent-based latency-aware adaptive resource allocation scheme. Simulation results are provided in Section III, and we conclude the paper in Section IV.
@@
II Intent-based Generative Semantic Multicasting
We consider a transmitter serving $K$ users in a multicasting setup as depicted in Fig. 1. The source signal to be transmitted to the users contains $L$ semantic classes. For example, assuming the source signal is images captured by a street scene camera, the semantic classes include “car”, “person”, “building”, “sky”, etc. Denoting the source signal by $\bf X$}, we have ${\bf X}=\bigcup_{l=1}^{L}{\bf X}_{l}$, where ${\bf X}_{l}$ represents the $l$’th class. Each user is only interested in a specific subset of the classes based on its goal or intent. For example, a user carrying out “traffic surveillance” based on the received signal is interested in the “car” class but not “sky”. This user may need details of the car class, e.g. the cars’ make and model, as well as their color and number plate. A user carrying out “weather monitoring” is only interested in details of the “sky” class, and another user carrying out “Path Planning” for Intelligent Transportation Systems (ITS) will be interested in the “street signs” and the “road” classes. Hence, we define the multi-user intent matrix ${\bf I}_{K\times L}$, where ${\bf I}]_{kl}=1$ if user $k$ is interested in class $l$, otherwise ${\bf I}]_{kl}=0$. The intended portion of the source signal for use $k$, i.e. the semantic classes each user is interested in, is thereby written as ${\bf X}^{(k)}_{I}=\{\bigcup_{l}{\bf X}_{l}|[{\bf I}]_{kl}=1\}$. Similarly, the non-intended portion of the source signal for use $k$ is thereby written as ${\bf X}^{(k)}_{N}=\{\bigcup_{l}{\bf X}_{l}|[{\bf I}]_{kl}=0\}$.
@@ II Intent-based Generative Semantic Multicasting
II-A Multi-class Semantic Decomposition for Mixed Reconstruction/Synthesis with Semantic Diffusion Models
II-A1 Multi-class Semantic Decomposition with DDRNets
@@
II-A2 Mixed Reconstruction/Synthesis with Semantic Diffusion Models
At the receiver, each user receives two datastreams, one containing its intended classes which is decompressed to reconstruct its intended portion of the source signal, and the other containing the semantic map $\bf M$} which is fed into a pre-trained generative diffusion model for high-fidelity synthesis of its non-intended portion. The signal retrieved at user $k$ is thereby written as ${\bf Y}^{(k)}={\bf Y}^{(k)}_{r}\bigcup{\bf Y}^{(k)}_{s}$, where ${\bf Y}^{(k)}_{r}$ is the reconstructed and ${\bf Y}^{(k)}_{s}$ is the synthesized portion. Any state-of-the-art generative models with semantic guidance can be used to locally synthesize the non-intended signal portions ${\bf Y}^{(k)}_{s}$ at the users. Here we pre-train GDMs with classifier-free guidance for high quality/fidelity image synthesis from the semantic map at each user. Note that we adopt diffusion models in the proposed framework as they enable synthesizing photorealistic outputs with better diversity, realism, and fine-grained details, thereby significantly improving users’ quality of experience, e.g. in comparison with other classes of generative models, e.g. GANs, VAEs, etc. We provide comparisons with GAN-based synthesis of the non-intended classes later in simulation results. Refer to [31, 32, 33, 34, 35] for details on diffusion models and their various recent applications in wireless communications.
Dropping the user index $k$ for brevity, the forward (diffusion) process is defined via
@@ II-A2 Mixed Reconstruction/Synthesis with Semantic Diffusion Models
where ${\bf Y}_{t}$ is the signal at the $t$’th diffusion step, $\rho^{\mathrm{Diff}}({\bf Y}_{t}\mid{\bf Y}_{t-1})$ is the diffusion distribution, and $\beta_{t}\in(0,1)$ is the trainable variance schedule controlling the amount of noise added at each step. To synthesize ${\bf Y}_{s}={\bf Y}_{0}$, the reverse (synthesis) process learns to remove the noise iteratively, transforming the noisy input ${\bf Y}_{\tau}$ to the synthesized image ${\bf Y}_{0}$, using the additional conditioning information from the received semantic map $^{\bf\hat{M}}$. The synthesis process is modeled by a parameterized Markov chain
[…]
@@
where $\epsilon$ is the true noise sampled from a Gaussian distribution, $\epsilon_{\theta}$ is the noise predicted by the model at time step $t$, given the noisy image ${\bf Y}_{t}$, the semantic map $^{\bf\hat{M}}$, and the time step $t$. During inference, the diffusion process iteratively denoises the noisy image input ${\bf Y}_{\tau}$ to generate the final image ${\bf Y}_{0}$. At each step $t$, the model predicts the noise $\epsilon_{\theta}({\bf Y}_{t},{\bf\hat{M}},t)$ and updates the image as
${\bf Y}_{t-1}=\frac{1}{\sqrt{1-\beta_{t}}}\left({\bf Y}_{t}-\frac{\beta_{t}}{\sqrt{1-\alpha_{t}}}\epsilon_{\theta}({\bf Y}_{t},{\bf\hat{M}},t)\right)+\sigma_{t}{\bf N},$ (4)
@@
II-B Intent-Aware Generative Multicasting